3(2x-7)+4-5x=2x(x-4)-(2x^2+3)
1
Expert's answer
2012-12-13T08:55:25-0500
3(2x-7)+4-5x=2x(x-4)-(2x^2+3)
6x-21+4-5x=2x^2-8x-2x^2-3
x-17=2x^2-2x^2-8x-3
x-17=-8x-3
x+8x=-3+17
9x=14
x=14/9